OpenStudy (anonymous):

A company's quality control department found an average of 5 defective modles for every 1000 models that were checked. If the company produced 60,000 models in a year, how many would expected to be defective? Please show work

OpenStudy (anonymous):

p=5/1000 n=60000 Ex=np

OpenStudy (anonymous):

dont understand

OpenStudy (anonymous):

you would do the probability of a model being defective, which is 5/1000, times by 60,000, which is 120 expected to be defective

OpenStudy (mathmate):

60000 * (5/1000)=60*5=300